# Joint Demonstration and Preference Learning Improves Policy Alignment with Human Feedback

### v2

@@ 6 Experiments
In this section, we present numerical results for the proposed method. More specifically, we intend to address the following questions: 1) How are the data efficiency and model improvement efficiency (w.r.t. the KL divergence violation) of our proposed method compared with other baselines? 2) Can we achieve the same convergence results after integrating the three stages into one stage? 3) And finally, how much does demonstration help in reward learning in AIHF? To answer these questions, we consider several applications of the alignment problem. In this section, we show the numerical results for large language model alignment in the Helpful and Harmless (HH) dataset (Bai et al., 2022) and robotics control problem in MuJoCo Todorov et al. (2012). In the appendix, we have another set of experiment where the goal is to align the language model so that it generates movie reviews with positive sentiment on IMDb dataset (Maas et al., 2011).
6.1 MuJoCo Tasks
@@
Figure 3: Left: Hopper Environment Middle: HalfCheetah Environment Right: Walker2d Environment; AIHF  (orange) vs RLHF(blue); results are averaged over 3 independent runs. We use 10k demonstrations and 20k preferences. The RLHF curve is initialized from a policy pre-trained by BC; the AIHF from a random policy. The performance is compared against the # of SAC steps performed (for AIHF each policy alignment performs 5k steps of SAC.)

### v3

@@ 5 Experiments
In this section, we provide numerical evaluations of the proposed method (5) (Alg. 1) and its variants (10) and (3.3), and comparing them with state-of-the-art methods RLHF (Ouyang et al., 2022), DPO (Rafailov et al., 2023), IPO (Calandriello et al., 2024) and SPIN (Chen et al., 2024). Our experiments demonstrate the advantages of the proposed methods in the following aspects: (1) Reward learning from demonstration and preference is the key to improving over standard RLHF. (2) Using demonstration in reward learning could increase model improvement efficiency (w.r.t. the KL divergence violation) (3) AIHF could reduce the effect of distribution mismatch caused by the sequential alignment method, thus breaking the performance limits of the state-of-the-art methods.
Models and datasets. In the first setting, we test Alg. 1 on Anthropic-HH (Bai et al., 2022) dataset [^2] Dataset available at https://huggingface.co/datasets/Anthropic/hh-rlhf. with (relatively small) Pythia (Biderman et al., 2023) models [^3] Models available at https://huggingface.co/EleutherAI. as policy models. Anthropic-HH is a preference dataset collected from 52B LLMs that provide two continuations based on helpfulness and harmlessness, and we pick 10k chosen/preferred continuation data to form the demonstration dataset, while others serve as preference dataset and RL prompt dataset. For the HH dataset, We first fine-tune the language models (Pythia-160M/1B/2.8B) through supervised fine-tuning over all chosen responses from the HH dataset for 1 epoch, we call it full-SFT model and use it as our base model. Moreover, we also SFT the language model using the selected top 10k chosen responses and name it as demonstration-SFT model. For each policy model, we use the exact same model Pythia-1.4B as the reward model.
The other setting we test is on 7B models. We use Ultrafeedback [^4] Available at https://huggingface.co/datasets/HuggingFaceH4/ultrafeedback_binarized. as our preference dataset (61.1k preference data) and Ultrachat200k [^5] Available at https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k. as the demonstration dataset (208k demonstration data), with mistral-7b-sft-beta  [^6] Available at https://huggingface.co/HuggingFaceH4/mistral-7b-sft-beta. (Jiang et al., 2023) as our base model. We use the same mistral-7b-sft-beta model as the initialization of the reward model.
Evaluation. For the Anthropic-HH dataset, we present the reward evaluated by the PKU-Alignment/beaver-7b-v3.0-reward model(Ji et al., 2024). In our 7B model experiments, we adopt the widely recognized HuggingFace Open LLM Leaderboard framework (Beeching et al., 2023). This evaluation suite measures LLM performance across six tasks: commonsense reasoning (Arc (Clark et al., 2018), HellaSwag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021)), multi-task language understanding (MMLU (Hendrycks et al., 2020)), mimicking human falsehoods (TruthfulQA (Lin et al., 2021)), and math problem-solving (GSM8K (Cobbe et al., 2021)). Additional implementation details can be found in Appendix A.1.
(a) Pythia-160M
(b) Pythia-1B
(c) Pythia-2.8B
Figure 2: Experiment results of Pythia-160M/1B/2.8B policy models, with the reward model trained from Pythia-1.4B. We record the average scores (across three trials) of AIHF and RLHF on the Anthropic-HH test dataset (See Tab. 2 in Appendix for more comparisons with other algorithms)
Results of small model (160M, 1B, and 2.8B) experiments. We observe that the proposed AIHF performs effectively when initiated from both the demonstration-SFT model and the full-SFT model. As shown in Fig.2, utilizing the same data, AIHF  algorithm can eventually outperform RLHF irrespective of the initial model. Furthermore, according to the numerical results as shown in Fig.3(a), we see that the proposed AIHF algorithm has a smaller deviation from the base model compared with the RLHF algorithm. This benefit of the AIHF approach is due to the fact that we incorporate the maximum likelihood IRL objective for both reward learning and policy learning. In this case, both the reward model and policy model will be trained to align with the demonstrations, which are also used in the training process of the SFT stage. We also conducted a study on the demonstration/preference data ratio in Fig. 3(b) and 3(c). We observe that AIHF consistently outperforms RLHF across different demonstration/preference data ratios. Furthermore, we also record the performance of the two variants, namely Direct AIHF (10) and Self-Play-AIHF (3.3), also RLHF, DPO, IPO, and SPIN in Tab. 2. Our proposed AIHF still outperforms all these methods in this experiment setting.
(a) KL divergence to demonstration-SFT policy
(b) AIHF vs RLHF with 10k demonstration, 5k preference
(c) AIHF vs RLHF with 10k demonstration, 10k preference
Figure 3: Experiment results on Pythia-1B policy models, where the reward model is trained from Pythia-1.4B models. We record the average scores of AIHF and RLHF on the Anthropic-HH test dataset, reporting the results across three different trials.
Results of large model (7B) experiments. We run AIHF, Direct AIHF, and Self-Play AIHF along with other methods on the 7B experiment setting. The results are presented in Fig. 4 (the numbers are recorded in Tab. 4), where we can see that similar to the 1B setting, AIHF is consistently outperforming other methods. Additionally, both Direct AIHF and Self-Play AIHF effectively outperform the RLHF model (zephyr-7b-beta). The success of AIHF, as well as Self-Play AIHF and Direct AIHF further suggests that joint learning from demonstration and preference is indeed beneficial for the alignment. We also conduct an ablation study with different choices of $w_{1}$ in (10), as shown in Tab. 6, the improvement of joint learning methods over baseline is robust. Furthermore, We also evaluate the reward models estimated using different methods (DPO, standard preference learning, and AIHF) over the widely used RewardBench (Lambert et al., 2024). The results, illustrated in Tab. 5 in the Appendix, show that the reward model trained through the AIHF can achieve significant improvement (especially on reasoning tasks) compared to both the standard BTL reward model in (3) and the implicit reward model in DPO.
Figure 4: Performance comparison between Direct AIHF, Self-Play AIHF training across the six benchmark datasets (See also Table 4 in the Appendix).
Other Results. Due to the page limits, we leave two additional experiments in the appendix: 1) movie review generation with positive sentiment on IMDb dataset (Maas et al., 2011), 2) experiment on Robotics control tasks in MuJoCo (Todorov et al., 2012). For the result of MuJoCo Experiment A.1.1, we observe that even though Behavior Cloning (BC)/SFT could provide a high-performing initialization, RLHF still fails to improve policy quality in the following RL stage. In the contract, AIHF can effectively integrate preferences and demonstrations, leading to a more robust reward function and consequently, a high-quality policy. For the IMDB result (Fig. 6), We show that AIHF is able to alleviate the distribution mismatch between the generated trajectories by the policy, and the data that the learned reward model is able to rank.
6 Conclusion
In this work, we study the alignment problem when diverse data sources from human feedback are available. Furthermore, we have developed an algorithmic framework that can integrate both expert demonstration and pairwise comparison data from human feedback to learn the reward functions for further guiding policy learning/model fine-tuning in the alignment pipeline. Through extensive evaluations on robotic control tasks and large language model alignment tasks, we demonstrate that our proposed method can outperform existing benchmarks on alignment tasks and is able to recover a better reward model to guide policy learning.
Appendix A Appendix
A.1 Experiment Setup and Additional Result
A.1.1 MuJoCo Tasks
@@
Figure 5: Top-Left: Hopper Environment; Top-Right: HalfCheetah Environment; Bottom: Walker2d Environment; AIHF  (orange) vs RLHF (blue) vs IPL (purple) (Hejna & Sadigh, 2024); results are averaged over 3 independent runs. We use 10k demonstrations and 20k preferences. The RLHF and IPL curve is initialized from a policy pre-trained by BC; the AIHF from a random policy. The performance is compared against the # of SAC steps performed (for AIHF each policy alignment performs 5k steps of SAC.)